2x Find a power series for the function f(x) and its interval of convergence. - 3.x2 5x – 2
Answer #1

\small \frac{{2x}}{{3{x^2} - 5x - 2}} = \frac{{2x}}{{(3x + 1)(x - 2)}} = \frac{A}{{x - 2}} + \frac{B}{{3x + 1}}\\ \\ A = \mathop {\lim }\limits_{x \to 2} \left( {(x - 2) \times \frac{{2x}}{{3{x^2} - 5x - 2}}} \right) = \mathop {\lim }\limits_{x \to 2} \left( {(x - 2) \times \frac{{2x}}{{(3x + 1)(x - 2)}}} \right)\\ \\ = \mathop {\lim }\limits_{x \to 2} \frac{{2x}}{{(3x + 1)}} = \frac{4}{7}

Also

\small B = \mathop {\lim }\limits_{x \to \frac{{ - 1}}{3}} \left( {\left( {x + \frac{1}{3}} \right) \times \frac{{2x}}{{3{x^2} - 5x - 2}}} \right)\\ \\ = \mathop {\lim }\limits_{x \to \frac{{ - 1}}{3}} \left( {\left( {x + \frac{1}{3}} \right) \times \frac{{2x}}{{(3x + 1)(x - 2)}}} \right)\\ \\ = \mathop {\lim }\limits_{x \to \frac{{ - 1}}{3}} \left( {\left( {x + \frac{1}{3}} \right) \times \frac{{2x}}{{3\left( {x + \frac{1}{3}} \right)(x - 2)}}} \right)\\ \\ = \mathop {\lim }\limits_{x \to \frac{{ - 1}}{3}} \frac{{2x}}{{3(x - 2)}} = \frac{{\frac{2}{3}}}{{3\left( {\frac{1}{3} + 2} \right)}} = \frac{2}{7}

That is \small \frac{{2x}}{{3{x^2} - 5x - 2}} = \frac{{2x}}{{(3x + 1)(x - 2)}} = \frac{4}{7}\frac{1}{{x - 2}} + \frac{2}{7}\frac{1}{{3x + 1}} .

Now \small \frac{1}{{x - 2}} = \left( { - \frac{1}{2}} \right)\left( {\frac{1}{{1 - \frac{x}{2}}}} \right) = \left( { - \frac{1}{2}} \right)\sum\limits_{n = 0}^\infty {{{\left( {\frac{x}{2}} \right)}^n}} = \sum\limits_{n = 0}^\infty {\left( {\frac{{ - 1}}{{{2^{n + 1}}}}} \right){x^n}} .....(i)

which is valid if \small {\left| {\frac{x}{2}} \right| < 1 \to |x| < 2} . That is converges if  \small x\in (-2,2).

Also \small \frac{1}{{3x + 1}} = \frac{1}{{1 - ( - 3x)}} = \sum\limits_{n = 0}^\infty {{{\left( { - 3x} \right)}^n}} = \sum\limits_{n = 0}^\infty {{{\left( { - 3} \right)}^n}{x^n}} .....(ii)

which is valid if \small |3x|<1\to |x|<\frac{1}{3} .Also interval of convergence is \small \left (-\frac{1}{3},\frac{1}{3} \right ) .

Now So both series valid is \small |3x|<1\to |x|<\frac{1}{3} . Also interval of convergence is \small \left (-\frac{1}{3},\frac{1}{3} \right ) .

Combinding (i) and (ii) we get \small \frac{4}{7}\frac{1}{{x - 2}} + \frac{2}{7}\frac{1}{{3x + 1}} = \frac{4}{7}\sum\limits_{n = 0}^\infty {\frac{{ - 1}}{{{{\left( 2 \right)}^{n + 1}}}}{x^n}} + \frac{2}{7}\sum\limits_{n = 0}^\infty {{{\left( { - 3} \right)}^n}{x^n}} = \sum\limits_{n = 0}^\infty {\frac{2}{7}\left( {\frac{{ - 1}}{{{{\left( 2 \right)}^n}}} + {{\left( { - 3} \right)}^n}} \right){x^n}} .

That is \small \frac{{2x}}{{3{x^2} - 5x - 2}} = \sum\limits_{n = 0}^\infty {\frac{2}{7}\left( {\frac{{ - 1}}{{{{\left( 2 \right)}^n}}} + {{\left( { - 3} \right)}^n}} \right){x^n}} valid if \small |x|<\frac{1}{3} . Also interval of convergence is \small \left (-\frac{1}{3},\frac{1}{3} \right ) .